OpenStudy (anonymous):

How do you find the nth term for a question like this? 1,8,27,64,125?

OpenStudy (anonymous):

they are the cube numbers:)

OpenStudy (anonymous):

looks like a sequence of cubes. what zara said

OpenStudy (anonymous):

\[ T_n=n^3 \]

OpenStudy (anonymous):

1=1*1*1 8=2*2*2 27=3*3*3 64=4*4*4 125=5*5*5

OpenStudy (anonymous):

Thank you so much everyone! I made that problem a whole lot harder than you guys made it. Thanks!

OpenStudy (anonymous):

Don't worry! I started off doing the wrong thing, I only realised half way through that they were cubes :')